An NBA Super Bowl player kicks a ball that spends 3s in the air and

lands 60 m ahead of him. Can you, the physicist, predict the speed at the top of the

trajectory without air resistance?
Answer #1

Horizontal distance moved by ball d = 60 m

total time of motion of ball t= 3s

speed along horizontal = d/t = 60 m /3s = 20 m/s

at top of motion there is only horizontal velocity and it remains constant so

at top speed of ball is 20 m/s.

thanks for asking.